What is boost parameters by Donald Marolf notes? explain in details


1
Expert's answer
2020-01-08T09:39:35-0500

According to Donald Marolf's Notes on Relativity and Cosmology (2003), "The boost parameter is another way to encode the information present in the relative velocity, and in particular it is a very natural way to do so from the viewpoint of Minkowskian geometry."

So what is a "boost"? Refer to the Notes:

"the transformation that changes reference frames is called a ‘boost.’ [...] Boosts are analogous to rotations.”

We simply use vector "r" and describe its position through the way it is rotated to some axis:


"x=r\text{ sin}\theta\\
y=r\text{ cos}\theta."


The slope is


"m=x/y=\text{ tan}\theta,"

and it is similar to

"v=x\/t."

Thus, different frames have different θ, which can be shown on any diagram.

"The parameter θ gives some notion of how different the two inertial frames (that of the moving observer and that of the stationary observer) actually are. For θ = 0, event A is at x = 0 and the two frames are the same, while for large θ event A is far up the hyperbola and the two frames are very different."

""\theta" is really a measure of the separation of the two reference frames. In this context, we also refer to θ as the boost parameter relating the two frames. The boost parameter is another way to encode the information present in the relative velocity"